General Info


This elog page is dedicated to the monitoring of tracker information during the DAQ shifts of 26 July to 4th August. One can update daily information on different operations undergone for the tracker during these shifts.


...


6 August 2010

The 515 (crate 5, power group 2) linear regulator problem:

  • Substituting TPSFE: spare TPSFE 95008f nothing changes! Additional problem with linear regulator S number 1 (actually there is a trip signal without failure!)
  • Substituting TDR: spare TDR 95077f replaces the 95084f TDR with the linear regulator K problem. The ploblem seems solved.
  • Putting back the TPSFE 95028f (since the other has this strange S linear regulator behaviour)
  • OK the 515 is back!

Pay attention to this "new" spares:

TPSFE 95008f Problem with a S linear regulator
TDR 95084f Problem with a K linear regulator


5 August 2010

17:20: TAS run with the tas_before.seq and tas_after.seq scripts (Matteo has to edit the config files ...). non-gaussian strips test:

0.1% (2) CalDB_1281022991.root 13358
1% (20) CalDB_1281022745.root 176
5% (102) CalDB_1281023118.root 140
10% (205) CalDB_1281022865.root 135

11:30: switched to RS442, updating the Lebedev libraries, ... everything OK.

09:30: Start Data Taking

  • Dynamic Pedestal: 1
  • No single channel exclusion
  • No gaussian channels
  • (requested occupancy table and sigma table: NEW CALIBRATION FORMAT)

Don't use the first two runs (several configuration problems)!


4 August 2010

19:00 From now:

  • No dynamic pedestals
  • No non-gaussian strip tagging
  • No single channel cluster exclusion

before 19:00 AMS flying out from the RAS!


3 August 2010

19:30 From now:

  • No dynamic pedestals
  • No non-gaussian strip tagging
  • Yes single channel cluster exclusion (n-side: 12 p-side: 10)

until midnight (occupancy tables not requested). Maybe this test has something wrong.

18:55 Non-gaussian strip test
Threshold Calibration N-Bad P-Bad
4 CalDB_1280858673.root 814 1358
8 CalDB_1280858673.root 306 465
16 CalDB_1280858488.root 153 217
32 CalDB_1280858096.root 69 96
64 CalDB_1280858239.root 63 74
128 CalDB_1280858375.root 63 69
512 CalDB_1280854850.root 48 48
1024 CalDB_1280854935.root 37 25

  • 15.20 - Installing DSP codes for TDRs and JINFs following the upgrades in the Kounine framework.
  • 13:37 - New Run Started (Run ID: 4C57FF76) with All Crates Powered ON. -> CalDB_1280835443.root
  • 12:19 - CalDB_1280830504 (12:20). This Calibration, done just before a TAS Run (old DSP code) showed no K-Side on Crate 5, TDR 15. Trying to understand the problem, a power failure on PWG3, Crate 5 was seen.
  • 12:22 - New Run Started (Run ID: 4C57EDB6) with New DSP code with Single channel cluster exclusion off and dynamic pedestals off and threshold for high occupancy channels (par 0xF) set to 0x800 (2048, that should means OFF). -> CalDB_1280830899.root
  • 09:28 - New Run Started (Run ID: 4C57C4E3) with 'yesterday's new DSP'. -> CalDB_*.root


2 August 2010

  • 23.40 - Switched on the single channel cluster exclusion with par=0x2828 (5 ADC for S and 5 ADC for K).
  • 23.10 - Switched on the single channel cluster exclusion with par=0xA0A0 (20 ADC for S and 20 ADC for K).
  • 22.50 - Switched on the single channel cluster exclusion with par=0x5060 (12 ADC for S and 10 ADC for K).
  • 21.45 - Single channel cluster exclusion off and dynamic pedestals off and threshold for high occupancy channels (par 0xF) set to 0x800 (2048, that should means OFF).
  • 21.00 - Single channel cluster exclusion off and dynamic pedestals off.
  • 18.30 - Data taking restarted with new new new (3fc3) DSP code with single channel cluster exclusion 0x5060 (12 ADC for S and 10 ADC for K) and dynamic pedestals (4).
  • PDS Power was switched OFF for PDS board replacement in Clean Room. NO power back until sometime in the afternoon.
  • 09:48 - The Run was Stopped in order to switch to AA Configuration..
  • 08:48 - New Run Started (Run ID: 4C566A48) with new DSP Code. -> CalDB_1280731717.root
    Here is a plot with the behaviour of the Event Size in the beginning (08:48) of a Run with the new DSP Code, as Matteo asked.

    EventSize_BeginningOfRunwithNewDSP_02_08_2010.png


1 August 2010

  • Note: The Event Size After the ~ 15:00 calibration WAS LOWER THAN BEFORE and before was TO HIGH. WE DON'T UNDERSTAND WHY...
  • 12:27 - New Run Started (RunID: 4C557118). New DSP coded was loaded successfully. -> CalDB_1280658411
  • 10:09 - New Run Started (RunID: 4C552B73). We failed to load the new DSP code (errors communicating with Crate 0) this morning so we are running with the old one. We will switch to the new one as soon as possible. -> CalDB_1280650096.root.
  • Not Yet but Have to Do: Starting the run (with the tracker temperature raising very fast) with the new DSP code, the single channel cluster exclusion with par=0x5060 (12 ADC for S and 10 ADC for K), and the dynamic pedestals (4) on.


31 July 2010

  • Playing with dynamic pedestals and single channel cluster exclusion.:

    EventSize_playing_with_0x1c_and_0xb_2.png
    EventSize_playing_with_0x1c_and_0xb_1.png
  • 23.25 - Single channel cluster exclusion to par=0xB0B0 (22 ADC for S and 22 ADC for K).
  • 23.05 - Single channel cluster exclusion to par=0x8080 (16 ADC for S and 16 ADC for K).
  • 22.50 - Switched on the single channel cluster exclusion with par=0x5060 (12 ADC for S and 10 ADC for K).
  • 22.30 - Switched off the single channel cluster exclusion and switched on the dynamic pedestals (8).
  • 21.50 - Switched on the single channel cluster exclusion with par=0x5060 (12 ADC for S and 10 ADC for K).
  • 21.25 - Switched off the single channel cluster exclusion and switched on the dynamic pedestals (4).
  • 20.20 - Problems with the JMDC so we need to restart it. This is the reason for the "hole" in the temperature monitor and after in the data taking...
  • 20.20 - Switched on the single channel cluster exclusion with par=0x5060 (12 ADC for S and 10 ADC for K).
  • 19.55 - Switched off the single channel cluster exclusion and switched on the dynamic pedestals (2).
  • 19.20 - Switched on the single channel cluster exclusion with par=0x5060 (12 ADC for S and 10 ADC for K).
  • 18.30 - Switched off the single channel cluster exclusion and switched on the dynamic pedestals (1).
  • 17.50 - New DSP code with dynamic pedestals off but new feature (single channel cluster exclusion) with par=0x5060 (12 ADC for S and 10 ADC for K).
  • EventSize at startup (tracker temperature increasing rapidly) with the old DSP code (official one) and dynamic pedestals ON. The first raising is probably due to silicon channels that are not yet marked as noisy (it were marked in the first next calibration), the raising of after is, instead, due to temperature so the dynamic pedestals are not fast enough. The steps, obviously are due to calibrations.:

    EventSize_at_startup_old_dsp.png

    * 10:19 - New Run Started (RunID: 4C53DC71). BB Configuration. Matteo's BPowerStartUp.sh Script was executed with no errors and artificial 200 musec artificial dead time setted for SDR. -> CalDB_1280564334.root


30 July 2010

  • EventSize with new DSP (3fb3) and changing, during the runs, the dynamic pedestals parameter (0, 1, 2, 4, 8). The step at 22.25 is due to a calibration.:

    EventSize_new_dsp_changing_dyn_ped_param.png
  • 23:45 - During the run dynamic pedestal were switched ON with the par to 8 (each event a channel pedestal is modified [increased or decreased] by 2 unit, so by 8/8 ADC counts). -> Event size increased very much.
  • 23:25 - During the run dynamic pedestal were switched ON with the par to 4 (each event a channel pedestal is modified [increased or decreased] by 2 unit, so by 4/8 ADC counts). -> Event size increased.
  • 23:00 - During the run dynamic pedestal were switched ON with the par to 2 (each event a channel pedestal is modified [increased or decreased] by 2 unit, so by 2/8 ADC counts). -> Event size increased a little.
  • 22.35 - During the run dynamic pedestal were switched ON with the par to 1 (each event a channel pedestal is modified [increased or decreased] by 1 unit, so by 1/8 ADC counts).
  • 22.00 - We modified the dead time to 200 musec. Now no more "ROOM" errors and the size is comparable to the previous version of new dsp code with dynamic pedestal off and with old dsp version with dynamic pedestals on.
  • 20.40 - After some problems with the BB configuration (we have the A0 Jinf-T0 disconnected...) we restarted and the new DSP code will be loaded (3fb3) and new configuration file. The run will be started with dynamic pedestals off. We' re running with 120 musec of dead time -> We have more "ROOM" errors and the size is a little bit bigger...
  • 15:43 - Run was Stopped and All SubD turned OFF as well as the PDS power (requested by clean room).
  • 14:32 - New Run Started (RunID: 4C52C676) after a first succesful TAS run for the day. -> CalDB_1280493171.root
  • 13:19 - New Run Started (RunID: 4C52B55C) -> CalDB_1280488793.root (13:22).
  • 12:45 - Problems with TAS Runs (unsolved for now).
  • 10:11 - New Run Started (RunID: 4C528909) after TAS Run was performed. -> CalDB_1280477446.root
  • 08:23 - New Run Started (RunID: 4C526FE4). All Tracker ON. Primary Cooling Pump ON. -> CalDB_1280471009.root (08:25). Note: when starting the run, the event size was huged! (> 3500).


29 July 2010

  • ~ 22:00 - Changed the set point to 25 degrees of TTCS to try to cool down the inner tracker. It worked but since we are without dynamic pedestal it increase the size. Set point came back to 26 degrees.
  • 17.30 - Looking at the event size the dynamic pedestals were disabled.
  • 16.30 - Moving to the new DSP code. - Probably with the dynamic pedestals on
  • 15.25 - Changing the configuration file for the old DSP (name of the conf file: 6001) in order to have the thresholds 3.5/1.0 automatically.
  • 15.25 - Loading the new DSP code, so at the next restart run we will have it running.
  • 15:05 - Run was Stopped and All Q-lists items disabled for communication cable operations in clean room.
  • 12:17 - New Run Started (RunID: 4C515511). Tracker has old DSP code threshold at 3.5/1.0. -> CalDB_1280398606.root
  • Several Starting-Up procedure tests.
  • 11:41 - Link Connexion with T0 was tested. A. Ruzkov disconnected A0 cable (connected to JINJ-1).
  • 10:50 - All SubD switched off and PDS was switched on clean room request. Primary pump switched off.
  • 09:11 - New Run Started (RunID: 4C512958). All Tracker ON. Primary Cooling Pump On. -> CalDB_1280387413.root OK. Old DSP code.


28 July 2010

  • During the evening/night the temperature of the tracker was a little bit higher and fluctuating. This has to be taken into account in the part of the run without the dynamic pedestals.
  • 21:58 - After the calibration (CalDB_1280346989.root) the size has come back to "normal".
  • 21:36 - After a TAS run, after re-loading the new DSP run, the dynamic pedestal were switched off. The size will decrease when we will calibrate...
  • 20:52 - Yes, the event size problem is with the dynamic pedestals...
  • 20:50 - Dynamic pedestal switched on again and let's see if the size will increase...
  • 19:26 - During the run the dynamic pedestals were switched off too see the behaviour in term of event size. After few minutes the size DECREASED very much and now it's COMPARABLE with the old DSP code...
  • 17:25 - Taking data again with the new DSP code. -> CalDB_1280330827.root
  • 17:20 - Coming back to old DSP code to perform TAS run -> CalDB_1280330570.root
  • 17:00 - We came back to the right status but we performed a TAS run with the new DSP code by mistake -> Obviously this TAS run is completely wrong! -> CalDB_1280329262.root
  • 16:30 - The calibration was not OK since the last operation left the system in a incoherent status -> CalDB_1280327257.root
  • 16:30 - Rozkov asked to check the AMSWire connection of the T0 and we switched off and on alla the T0.
  • ~ 16:00 - Starting with the new DSP code (3.5/1.0) -> CalDB_1280325365.root
  • 15:16 - New Run Started (RunID: 4c502db8) with thresholds at 3.5/1.0. CalDB_1280322886.root
  • 14:50 - DAQ Procedure Tests. Power was shut down on all crates.
  • 14:15 - New Run Started (RunID 4C501F16) with thresholds at 3.5/1.0.
  • 10:15 to ~14:00, threshold back at 4.0/2.0 (because of initialization of DAQ procedures done around 10:15)
  • 10:31 - DAQ Stopped, X. Cai testing procedures. Tracker powered off at some point, temporarily. New Q-lists names which issued in no temperature monitoring during ~1-2 hour.
  • 10:15 - TTCS tests during ~30min.
  • 09:12 - CalDB_1280301066 - Fist Calibration OK.
  • 09:20 - Primary TTCS pump was correctly switched on.
  • 09:00 - All Crates Powered On. Run Started, RunID: 4C4FD801. Running with thresholds at 3.5/1.0. Probably these thresholds weren't really applied. Check them!


27 July 2010

  • 01:00 - We stop the run and switched off the detectors. We mantained on the TTCE and M-Crate temperature monitor.
  • 23:30 - We installed new DSP code (3fa3) with new configuration file (7011) and we started to take data with thresholds at 3.5/1.0
  • During the afternoon we changed from 1up-1down to any1 and then we come back to 1up-1down.
  • 22:00 - The bug of above was fixed by Alberto and now we can look at all the 192 cabled ladders.
  • 20:30 - Found a probable bug in MonitorUI that shows the 4 ladder decabled (obviously with empty data!) and probably not 4 of the ladders in the plane 9N.
  • 18:30 - Running with thresholds at 3.5/2.0.


26 July 2010

  • Monitoring of Tracker Temperatures with all SubD powered on. Attention carried to Plane 1N temperatures (TRD on and MLI)
  • 18:00 - Failure on JMDC, no more monitoring of temperatures during 2 hours but all SubD remained powered on.

-- PierreSaouter - 27-Jul-2010

-- AlbertoOliva - 06-Aug-2010

Topic attachments
I Attachment History Action Size Date Who Comment
PNGpng EventSize_BeginningOfRunwithNewDSP_02_08_2010.png r1 manage 22.6 K 2010-08-02 - 10:08 PierreSaouter Event Size with New DSP at beginning of Run
PNGpng EventSize_at_startup_old_dsp.png r1 manage 28.1 K 2010-07-31 - 11:30 PierreSaouter EventSize at startup (tracker temperature increasing rapidly) with the old DSP code (officvial one) and dynamic pedestals ON. The dynamic pedestals are not fast enough...
PNGpng EventSize_new_dsp_changing_dyn_ped_param.png r1 manage 29.1 K 2010-07-31 - 11:27 PierreSaouter EventSize with new DSP (3fb3) and changing, during the runs, the dynamic pedestals parameter (0, 1, 2, 4, 8)
PNGpng EventSize_playing_with_0x1c_and_0xb_1.png r1 manage 29.5 K 2010-07-31 - 23:42 PierreSaouter Playing with dynamic pedestals and single channel cluster exclusion.
PNGpng EventSize_playing_with_0x1c_and_0xb_2.png r2 r1 manage 29.9 K 2010-07-31 - 23:51 PierreSaouter Playing with dynamic pedestals and single channel cluster exclusion.
Texttxt TRACKER10_ams.txt r1 manage 0.4 K 2010-08-07 - 19:12 SadakazuHaino Beam position and angle (x/y/z/dx/dy/dz) in AMS coord.
Texttxt TRACKER280_ams.txt r1 manage 11.5 K 2010-08-07 - 19:16 SadakazuHaino Beam position and angle (x/y/z/dx/dy/dz) in AMS coord.
Texttxt TRACKER416_ams.txt r1 manage 17.1 K 2010-08-07 - 19:14 SadakazuHaino Beam position and angle (x/y/z/dx/dy/dz) in AMS coord.
Texttxt TRACKER60_ams.txt r1 manage 2.5 K 2010-08-07 - 19:12 SadakazuHaino Beam position and angle (x/y/z/dx/dy/dz) in AMS coord.
Texttxt TRACKER80_ams.txt r1 manage 3.3 K 2010-08-07 - 19:15 SadakazuHaino Beam position and angle (x/y/z/dx/dy/dz) in AMS coord.
Edit | Attach | Watch | Print version | History: r49 < r48 < r47 < r46 < r45 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r49 - 2010-09-30 - AlbertoOliva
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback